# WikiVideo: Article Generation from Multiple Videos

### v1

@@ 5.1 Experiment 1: CAG and Baselines
Table 2 shows WikiVideo article generation results comparing CAG and the baselines described above. We find that simple concatenations of the per-video summaries—whether the initial generic ones ($\textsc{Concat}_{\textsc{Gen}}$) or those obtained via reprompting ($\textsc{Concat}_{\textsc{RePrompt}}$)—yield articles of very poor quality. Although manual inspection reveals these per-video summaries to contain mostly accurate descriptions of scenes and notable visual entities (e.g. the Eiffel tower), we take this as compelling evidence that individual video summaries are thoroughly inadequate for our task, absent higher-level synthesis.
@@
Results with CAG-0 and CAG, both of which incorporate the text-only aggregator LLM, offer further evidence for this interpretation, as we observe large gains across all metrics for both of these methods relative to the Concat baselines. For most metrics, CAG also obtains superior results to CAG-0, suggesting that supplying the aggregator with the RePrompted summaries (in addition to the generic ones) further enhances article quality.
Table 3 provides a qualitative example comparing a summary for a single video about the 2019 Notre Dame Cathedral fire generated by QwenVL2.5 (left) and the final article output by QwenVL2.5-based CAG (right). While the single video summary correctly identifies the location (Paris) and the most salient visual entity (Notre Dame), the descriptions are alternately too granular (e.g. the surrounding buildings are trees are overly obscured by smoke) and too florid (e.g. the scene is both dramatic and tragic, capturing a moment of significant destruction to a beloved landmark) for outputs intended to be read by humans, like a Wikipedia-style article on the event. By contrast, the CAG-generated article achieves a much more appropriate level of granularity, although generation of ungrounded details (bolded) remains a challenge—despite our prompt excluding the target event mention and explicitly prohibiting reliance on non-video knowledge sources.
Method VideoLLM R1 R2 RL BS Arg AS
$\textsc{Concat}_{\textsc{Gen+audio}}$ LLaVA-Video 5.21 1.30 3.54 79.49 20.25 5.95
VAST 16.88 1.59 11.78 80.44 8.20 8.20
InternVideo2.5 11.56 2.33 7.83 80.72 18.72 8.48
QwenVL2.5 11.05 3.20 6.89 81.63 22.11 7.50
[…]
InternVideo2 32.79 8.82 19.18 85.81 24.77 13.13
QwenVL2.5 32.05 9.17 18.99 85.74 26.25 12.62
CAG-2 QwenVL2.5 33.96 10.90 19.45 86.35 30.80 14.29
@@
Table 4: WikiVideo article generation results for CAG and baselines with audio inputs. Bottom row shows CAG without audio (copied from Table 2), which performs best.
5.2 Experiment 2: Incorporating Audio
Articles in WikiVideo have many claims grounded partly or only in videos’ audio signal  (Table 1). Here, we consider the impact of adding audio information as additional input to CAG and to $\textsc{Concat}_{\textsc{Gen}}$. For all methods except for VAST–which takes raw audio input–we transcribe the audio of each video using whisper-v3 large (Radford et al., 2022). For the $\textsc{Concat}_{\textsc{Gen}}$ baseline, we provide the transcription as additional input alongside the instruction and frames to the VideoLLM. For CAG, we provide the transcriptions for each video together with the per-video summaries as input to the text-only aggregator LLM.
@@ 5.2 Experiment 2: Incorporating Audio
Table 4 presents the results. Similar to the previous experiment, we find that $\textsc{Concat}_{\textsc{Gen}}$ continues to yield poor quality articles, even with audio information. Notably, however, both $\textsc{Concat}_{\textsc{Gen}}$ and CAG consistently obtain worse results with audio inputs than without (Table 2)—despite the sizable fraction of audio-support claims in WikiVideo. For $\textsc{Concat}_{\textsc{Gen}}$, this may partly be explained by the fact that the pretraining data for the VideoLLMs we study does not include audio transcripts, and thus the prompts that incorporate them are out-of-distribution. For CAG, we found that including audio transcripts tended to result in substantially shorter final articles (avg. $\sim 164$ tokens) than omitting them (avg. $\sim 206$ tokens)—suggesting that the former may be less thorough in their coverage of the event relative to the references, leading to lower metric scores (Appendix E has quantitative examples). Identifying ways to more effectively incorporate audio into the WikiVideo task thus constitutes an intriguing direction for future work.
Method Retriever VideoLLM R1 R2 RL BS Arg AS
CAG-2 V-ColBERT IV 20.46 3.83 12.77 82.74 17.24 7.49
QVL 24.13 4.68 14.42 83.67 20.93 10.59
MMMORRF IV 20.52 3.10 12.63 76.61 16.71 8.84
QVL 23.84 4.91 14.32 77.85 20.65 9.01
Oracle QVL 33.96 10.90 19.45 86.35 30.80 14.29
@@
Table 5: Results with CAG using different retrievers. The top 5 videos in a ranked list are used for generation. V-ColBERT=Video-ColBERT; IV=InternVideo2.5; QVL=QwenVL2.5. Oracle retrieval results are from Table 2.
5.3 Experiment 3: Retrieval Augmented Generation
@@
In contrast to the previous two experiments, which were run using only relevant videos for each target event (the oracle setting), here we consider the RAG setting, in which relevant videos must be retrieved. We use the full set of MultiVENT 2.0 (Kriz et al., 2024) videos from the test set as our corpus (109K videos).
We perform retrieval under two settings: video-only, and audio-visual. In the video-only setting, we create our index using VideoColBERT (Reddy et al., 2025), an efficient bi-encoder method for token-wise video retrieval. We create 26 vectors to represent the spatial (12) and spatio-temporal (14) features and train on the subset of videos in the MultiVENT 2.0 training split that have queries associated with them. In the audio-visual setting, we use all features from the videos to create the index, which includes visual frame features, extracted OCR, and audio transcripts. We use MMMORRF (Samuel et al., 2025), the state-of-the-art retrieval method on MultiVENT 2.0, which combines these features to produce ranked lists. We generate articles using the top 5 videos for each query.
@@ 6 Conclusion
In this paper we introduce the difficult task of automatically generating Wikipedia-style articles based on multiple videos about real-world events. To facilitate work on this task, we collect and release WikiVideo, a benchmark of high-quality, expert-written articles grounded in diverse videos, ranging from amateur footage to professional news coverage, which are densely annotated for multimodal support of the articles’ claims. Further, since existing systems for video-based summarization tasks are memory-intensive and overly focused on low-level video descriptions, we introduce Collaborative Article Generation (CAG)—a strong baseline for our task that leverages elements of relevance feedback and test-time scaling to iteratively construct high-level event-centric summaries. Our experiments demonstrate the effectiveness of CAG compared to alternative baselines—both in oracle retrieval and RAG settings. While CAG takes a significant step toward addressing the above limitations, other challenges remain—such as the need for improved methods for integrating audio signal into the article generation process and more effective training of VideoLLMs to better support this task—which represent exciting directions for future work.
7 Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE2139757. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. We thank Neha Verma and Marc Marone for their feedback on writing and narrative.
@@ Appendix A Dataset Statistics
(b) Video Statistics
Table 6: WikiVideo Statistics
6(a) contains additional statistics about the claims in WikiVideo and 6(b) has additional statistics about the videos. Further examples of WikiVideo articles can be found in Table 18 and Table 19. We also report the IAA metrics in Table 7.
@@ Appendix B Data Collection
This appendix discusses the annotation process for WikiVideo in greater detail. Beyond topic selection, recall that the annotation process includes claim decomposition (subsection B.1), subclaim rewriting (subsection B.2), subclaim grounding (subsection B.3), and article rewriting (subsection B.4).
@@ B.1 Claim Decomposition
Much recent work has studied the appropriate granularity of subclaims in a claim decomposition and the applications of such decompositions to natural language inference (NLI) and claim verification (Min et al., 2023; Gunjal & Durrett, 2024; Wanner et al., 2024b; a; Hu et al., 2025; Srikanth & Rudinger, 2025). Our own claim decomposition method is most similar to those of Gunjal & Durrett (2024) and Wanner et al. (2024a) in that we decontextualize subclaims—insert elided or abstracted context (e.g. by substituting pronouns with named entities)—up to the point that no further verification of extracted facts is required. (An example of one of our claim decomposition prompts can be found in Figure 4). Although not a rigourous form of decomposition, this notion is straightforward to apply during annotation. Additionally for simplicity, we treat dates as named entities and do not decompose them beyond their mention as given in the text. For example, some methods, like Wanner et al. (2024b), may decompose the claim “The event occurred on 15 April 2019” as:
@@ Appendix C Relevance and RePrompt
In this section, we include the prompt used with the reasoning model (Figure 10) and examples of real queries provided to the VideoLLM (Figure 11). We note that we had to add force the model to produce the prefix ”Describe the video in detail and focus on …” or else sometimes the reasoner would not follow our instructions and produce a new query.
@@ Appendix E Qualitative Differences Between Outputs
Appendix F Additional Results

### v2

@@ 5.1 CAG and Baselines
Table 2 shows WikiVideo article generation results comparing CAG-R against the baselines described above. We find that simple concatenations of the per-video summaries—whether the initial generic ones (Concat-0) or those obtained via reprompting (Concat-RePrompt, Concat-R)—yield articles of poor quality. Although manual inspection reveals these per-video summaries to contain mostly accurate descriptions of scenes and notable visual entities (e.g. the Eiffel tower), we take this as compelling evidence that individual video summaries are inadequate for our task, absent higher-level synthesis.
@@
Results with CAG-0 and CAG-R, both of which incorporate the text-only aggregator LLM, offer further evidence for this interpretation, as we observe large gains across all metrics for both of these methods relative to the Concat baselines. For most metrics, CAG-R also obtains superior results to CAG-0, suggesting that supplying the aggregator with the reprompted summaries (in addition to the generic ones) further enhances article quality.
5.2 Generalization to Unseen Events
LLMs are trained on vast quantities of internet text, including Wikipedia Barham et al. (2023); Soldaini et al. (2024); Cheng et al. (2024). As such, the LLMs underlying each VLM in our experiments can be assumed to have parametric knowledge about WikiVideo events from 2016 to 2024. To illustrate the generalizability of CAG beyond these events, Table 3 reports results on a subset of WikiVideo containing events that occurred after the latest knowledge cutoff date of any of these models. [^7] Qwen2.5 Release: 2024, Earliest 2025 Event: January 15. Here, we find that CAG obtains superior results to the Qwen 2.5 LLM, indicating that it exhibits better generalization to genuinely novel events.
Method R1 BS Arg AS G
QVL 15.6 82.6 25.6 15.9 -
CAG 42.0 85.7 32.8 14.0 51.2
LLM 33.5 84.5 22.9 18.4 19.5
Table 3: Results on WikiVideo-25 (only from 2025). LLM- Qwen 2.5 LLM. G: human annotated grounding.
This experiment also highlights an interesting finding: VLMs struggle to connect visual signals to parametric knowledge. This can be see in Table 3, where the performance of Concat-0 is similar for all events and the 2025 subset. Appendix E shows summaries for a single video about the 2019 Notre Dame Fire generated by QwenVL2.5 (left) and CAG (right). While the single video summary correctly identifies the location (Paris), the most salient entity (Notre Dame), and the physical event (fire), the output doesn’t make any high-level inference about the event and it fails to connect the visual information to the event in parametric knowledge. We suggest two possible reasons for this: (1) the backbone LLM loses the ability to perform high-level inference because text-video pretraining data is purely focused on low-level descriptions Zhang et al. (2024a), or (2) the LLM does not have the ability to activate this parametric knowledge from the projected visual signals.
5.3 Grounding in Videos
Information presented in the generated articles should be fully grounded in the videos provided as context. To evaluate this, we have human annotators perform the WikiVideo claim grounding annotation task on the model predictions from Table 3 for CAG and the Qwen2.5 LLM. We score the Groundedness of an article as the mean number of claims judged as supported by the video content: $G=\frac{1}{|C|}\sum_{c_{i}\in C}f(c_{i},V)$, where $C$ is the set of claims in the output, $V$ is the set of associated videos, and $f$ is scored as:
$f(c,V)=\begin{cases}1&\text{$c$ is supported by some $v_{i}\in V$}\\ 0&\text{otherwise}\end{cases}$ (1)
In Table 3, we find that a majority of the claims from CAG are grounded in the article—a stronger performance than the LLM-only generated responses. However, the difference between Rouge and BertScore does not reflect the poor quality of LLM generations or the lack of grounding, highlighting a need for strong evaluation metrics for video-to-text tasks. [^8] Appendix E for qualitative comparison between CAG and LLM.
5.4 Experiment 5: Including Audio
Method VLM R1 BS Arg AS
CAT-0 QVL 11.1 81.6 22.1 7.50
CAG-R QVL 32.1 85.7 26.3 12.6
CAG-R QVL 34.0 86.4 30.8 14.3
@@ 5.4 Experiment 5: Including Audio
Table 4: WikiVideo article generation results for CAG with audio inputs. Bottom row shows CAG without audio (copied from Table 2), which performs best.
@@
Articles in WikiVideo have many claims grounded partly or only in videos’ audio signal  (Table 1). Here, we consider the impact of adding audio information as additional input to CAG. In Table 16 we observe consistently worse results with audio inputs than without. We found that including audio transcripts tended to result in substantially shorter final articles (avg. $\sim 164$ tokens vs. $\sim 206$ tokens)—suggesting that they may be less thorough in their coverage of the event relative to the references. Identifying ways to more effectively incorporate audio into the WikiVideo task is thus an intriguing direction for future work. [^9] A longer discussion of audio including is in Appendix F.
Retriever VLM R1 BS Arg AS
VC QVL 24.1 83.7 20.9 10.6
MRF QVL 23.8 77.9 20.7 9.0
Oracle QVL 34.0 86.4 30.8 14.3
@@
Table 5: Results with CAG using different retrievers. The top 5 videos in a ranked list are used for generation. Oracle retrieval results are from Table 2. VC: Video-ColBERT, MRF: MMMORRF
5.5 Experiment 5: Retrieval Augmented Generation
@@
In contrast to the previous two experiments, which were run using only relevant videos for each target event (the oracle setting), here we consider the RAG setting, in which relevant videos must be retrieved. We use the full set of MultiVENT 2.0 (Kriz et al., 2024) videos from the test set as our corpus (109K videos). We perform retrieval from this collection of videos with two retrievers: Video-ColBERT (VC in results; Reddy et al., 2025) and MMMORRF (MRF; Samuel et al., 2025), the state-of-the-art retrieval method on MultiVENT 2.0. We generate articles using the top 5 videos.
@@ 6 Conclusion
In this paper we introduce the difficult task of automatically generating Wikipedia-style articles based on multiple videos about real-world events. We collect and release WikiVideo, a benchmark of high-quality, expert-written articles grounded in diverse videos, ranging from amateur footage to professional news coverage, which are densely annotated for multimodal support of the articles’ claims. Further, since existing systems for video-based summarization tasks are memory-intensive and overly focused on low-level video descriptions, we introduce Collaborative Article Generation (CAG)—a strong baseline for our task that leverages elements of relevance feedback and test-time scaling to iteratively construct high-level event-centric summaries. Our experiments demonstrate the effectiveness of CAG compared to alternative baselines—both in oracle retrieval and RAG settings. While CAG takes a significant step toward addressing the above limitations, future work remains in: (1) efficient multi-video inference (single inference step), (2) effectively integrating audio signal into the article generation process, (3) training VLMs to perform high-level inference, (4) connecting visual signals to parametric knowledge, and (5) improving video retrieval for RAG performance.
Limitations
Parametric Knowledge
Models trained on Wikipedia Barham et al. (2023); Cheng et al. (2024) will be able to artificially inflate scores as quoting from a Wikipedia article leads to high scores across all metrics than human annotators (see Appendix H). We suggest reporting results on each subset (WikiVideo-24 and WikiVideo-25) to help with this limitation. However, when new models come out trained on Wikipedia articles from 2025, the data will again need to be updated.
Computational Costs
Multi-video inference is an important problem to solve and a limiting factor of our method. To run inference on the 72B version of CAG, it requires 8 80GB A100s and this is only for single video inference. For multi-video this will continue to compound the computational costs and future work should be dedicated to decreasing these costs.
Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE2139757. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Dataset Statistics
Table 6 contains additional statistics about the claims in WikiVideo and Table 7 has additional statistics about the videos. Further examples of WikiVideo articles can be found in Table 21 and Table 22. We also report the IAA metrics in Table 8.
@@
Table 7: Video Statistics
@@ Appendix B Data Collection
This appendix discusses the annotation process for WikiVideo in greater detail. Beyond topic selection, recall that the annotation process includes claim decomposition, subclaim rewriting, subclaim grounding, and article rewriting.
@@ B.1 Claim Decomposition
Much recent work has studied the appropriate granularity of subclaims in a claim decomposition and the applications of such decompositions to natural language inference (NLI) and claim verification (Min et al., 2023; Gunjal and Durrett, 2024; Wanner et al., 2024b, a; Hu et al., 2025; Srikanth and Rudinger, 2025). Our own claim decomposition method is most similar to those of Gunjal and Durrett (2024) and Wanner et al. (2024a) in that we decontextualize subclaims—insert elided or abstracted context (e.g. by substituting pronouns with named entities)—up to the point that no further verification of extracted facts is required. (An example of one of our claim decomposition prompts can be found in Figure 4). Although not a rigourous form of decomposition, this notion is straightforward to apply during annotation. Additionally for simplicity, we treat dates as named entities and do not decompose them beyond their mention as given in the text. For example, some methods, like Wanner et al. (2024b), may decompose the claim “The event occurred on 15 April 2019” as:
@@ Appendix C Relevance and RePrompt
In this section, we include the prompt used with the reasoning model (Figure 10) and examples of real queries provided to the VideoLLM (Figure 11). We note that we had to add force the model to produce the prefix "Describe the video in detail and focus on …" or else sometimes the reasoner would not follow our instructions and produce a new query.
@@ Appendix E Qualitative Differences Between Outputs
Qualitative Results for 2025 Subset
We also present the qualitative differences between CAG and LLM responses for the 2025 subset of WikiVideo. The LLM is prompted with "Generate a Wikipedia article about [event]." Table 15 shows the comparison in these results. However, the LLM results are poor, filled with ungrounded and not factual information. In the main text, it’s noted that the Rouge and BertScores show similar performance of the LLM to CAG-R. However, these qualitative examples show that the LLM is only matching the style of Wikipedia articles and not generating meaningful and factual text.
Topic LLM CAG
Liberation Day Tariffs Liberation Day Tariffs, also known as Vrijheidsdag Tariffs, are special rates and discounts offered by various businesses and organizations in the Netherlands on May 5th, Liberation Day (Vrijheidsdag). This day commemorates the end of the Nazi occupation during World War II in 1945. The tariffs typically include reduced prices on public transport, free admission to museums, and special events and activities designed to encourage participation and celebration among the Dutch population. These offers aim to make the commemoration accessible and inclusive, fostering a sense of community and national pride. On April 2, 2025, President Donald Trump delivered a speech in Washington, D.C., announcing new tariffs on imports from several countries, including China, the European Union, Vietnam, Taiwan, Japan, India, South Korea, Thailand, Switzerland, Indonesia, and Malaysia. The tariffs, ranging from 20% to 36%, were designed to address what the administration perceived as unfair trade practices and to encourage domestic manufacturing. Trump emphasized that the U.S. would not impose a full 100% reciprocal tariff, stating his intention to be k̈ind to the worldẅhile protecting American industries. The announcement, made after the stock market closing bell, had already contributed to market volatility, with significant losses reported in the stock market, including a \$3 trillion drop in value, referred to as a s̈tock market armageddon.F̈inancial analysts, such as those from JPMorgan Chase & Co., raised the recession risk forecast to 60%, up from 40% prior to the announcement. The legal and economic implications of these tariffs were widely discussed, with experts debating the constitutionality of the President’s actions under the International Emergency Economic Powers Act (IEEPA) and the potential long-term impacts on global trade dynamics and consumer prices.
Blue Ghost Mission 1 The Blue Ghost Mission 1 was the inaugural flight of the Blue Ghost spacecraft, developed by Astra, an American aerospace manufacturer. Launched from the Mid-Atlantic Regional Spaceport (MARS) on Wallops Island, Virginia, the mission aimed to demonstrate the capabilities of the reusable suborbital vehicle designed for research and technology development. The spacecraft, which can carry payloads up to 150 kilograms, successfully reached its intended altitude and completed a series of tests before safely returning to Earth. The mission marked a significant milestone in Astra’s efforts to provide cost-effective and flexible access to space for scientific and commercial applications. The Blue Ghost M1 lunar lander, developed by Firefly Aerospace, successfully landed on the Moon on February 24, 2025, as part of NASA’s Artemis program. The mission, which aimed to advance lunar exploration and scientific research, was a significant milestone in space exploration. The spacecraft, equipped with advanced scientific instruments, was designed and assembled at Firefly Aerospace’s industrial facility, where engineers and technicians worked meticulously to ensure its precision and reliability. The landing process was closely monitored from the company’s control room, where a team of experts tracked the lander’s descent using detailed telemetry data. The successful touchdown was met with jubilation among the team members, marking a testament to their dedication and innovation in space technology. The mission not only contributed to the ongoing exploration of the Moon but also highlighted the capabilities of private aerospace companies in collaboration with government space agencies.
Canadian Federal Election The 2025 Canadian federal election was held on October 21, 2025, to elect members of the House of Commons of Canada, the lower house of the bicameral Parliament of Canada. This election determined the composition of the 45th Canadian Parliament and the next Prime Minister of Canada. The Liberal Party, led by incumbent Prime Minister Justin Trudeau, faced off against the Conservative Party, led by Pierre Poilievre, among other smaller parties. Key issues included economic management, healthcare reform, and climate change policies. The election saw high voter turnout and significant debate over the direction of national policy, reflecting deep divisions within the electorate. Results from the election were closely watched, with implications for both domestic and international affairs. The 2025 Canadian federal election, held on April 29, 2025, resulted in a fragmented Parliament with no single party achieving a majority. The Liberal Party, led by the incumbent Prime Minister, secured 119 seats, while the Conservative Party, under the leadership of Pierre Poilievre, gained over 20 seats, bringing their total to 83 and receiving the highest share of votes since 1988. The Bloc Qu0̆0e9b0̆0e9cois won 15 seats, the New Democratic Party (NDP) secured only 4 seats, and the Green Party failed to win any seats. The election saw significant shifts in voter support, particularly from the NDP to the Liberals, and created a challenging political environment with neither the NDP nor the Liberals able to form a coalition government. Pierre Poilievre, in his post-election speech, acknowledged the party’s progress but emphasized the need for continued effort and change, looking ahead to future elections. The overall atmosphere during election night was one of excitement and celebration, especially among Liberal supporters, as they awaited the final results.
Table 15: Qualitative outputs on a sample of the 2025 subset of WikiVideo for the LLM and CAG outputs.
Appendix F Incorporating Audio
Method VLM R1 BS Arg AS
CAT-0 QVL 11.1 81.6 22.1 7.50
CAG-R QVL 32.1 85.7 26.3 12.6
CAG-R QVL 34.0 86.4 30.8 14.3
Table 16: WikiVideo article generation results for CAG with audio inputs. Bottom row shows CAG without audio (copied from Table 2), which performs best.
Articles in WikiVideo have many claims grounded partly or only in videos’ audio signal (Table 1). Here, we consider the impact of adding audio information as additional input to CAG and to Concat-0. For all methods except for VAST–which takes raw audio input–we transcribe the audio of each video using whisper-v3 large (Radford et al., 2022). For the Concat-0 baseline, we provide the transcription as additional input alongside the instruction and frames to the VideoLLM. For CAG, we provide the transcriptions for each video together with the per-video summaries as input to the text-only aggregator LLM.
Table 16 presents the results. Similar to the previous experiment, we find that Concat-0 continues to yield poor quality articles, even with audio information. Notably, however, both Concat-0 and CAG consistently obtain worse results with audio inputs than without (Table 2)—despite the sizable fraction of audio-support claims in WikiVideo. For Concat-0, this may partly be explained by the fact that the pretraining data for the VideoLLMs we study does not include audio transcripts, and thus the prompts that incorporate them are out-of-distribution. For CAG, we found that including audio transcripts tended to result in substantially shorter final articles (avg. $\sim 164$ tokens) than omitting them (avg. $\sim 206$ tokens)—suggesting that the former may be less thorough in their coverage of the event relative to the references, leading to lower metric scores (Appendix E has quantitative examples). Identifying ways to more effectively incorporate audio into the WikiVideo task thus constitutes an intriguing direction for future work.
Appendix G Additional Results